OpenStudy (anonymous):

write the equation of the line through -1 and -2 that is parallel to 3x+ 4y=12 please tell me the right answer

OpenStudy (anonymous):

First, find the slope of the line (1): y = -3x/4 + 3. Write the equation of the line (2), parallel to line (1), and passes through point (-1, -2): y = -3x/4 + b . Replace x by (-1) and y by (-2). (-2) = -3(-1)/4 + b b = -8/3. Finally the answer is : y = .....